### v1

@@ 2.5 The demand and supply shocks in the global oil market affect the US economy in different ways
Figure 5 shows how the level of US real GDP and the level of the CPI respond to the structural shocks. The solid, grey lines show impulse responses estimated using the original data and the replication code. The associated two-standard-error bands are shaded in green. The impulse responses estimated using Kilian’s (2009a) code agree exactly, which produces a striped effect. The dotted, blue lines show the upper and lower ends of the confidence intervals using Kilian’s (2009a) code. These estimates produce slightly narrower confidence intervals, likely due to the fact that the block length used to produce the shaded region is two quarters longer. Nevertheless, the difference is minor. Figure 5 is directly comparable to Kilian’s (2009a) figure 5.
@@
Note: Clockwise from top left, cumulative response of real GDP to an unanticipated disruption in crude-oil supply; cumulative response of the price level to an unanticipated disruption in crude-oil supply; cumulative response of the price level to an unanticipated increase in aggregate demand; cumulative response of the price level to an unanticipated increase in oil-specific demand; cumulative response of real GDP to an unanticipated increase in oil-specific demand; cumulative response of real GDP to an unanticipated increase in aggregate demand. Solid lines are based on the point estimates from the model in (4). Shaded areas show two-standard-error bands constructed using a block bootstrap procedure. Dotted lines show the point estimates and two-standard error bands produced by Kilian’s (2009a) code. All statistics in the figure are based on the original data (Kilian, 2009b).
Figure 6 updates figure 5. To do this, data from February 1974 to March 2024 are first used to estimate the structural VAR model. Then the associated monthly series of structural shocks are averaged by quarter to produce the quarterly averages used in the second-stage regressions in (4).
@@
Note: Clockwise from top left, cumulative response of real GDP to an unanticipated disruption in crude-oil supply; cumulative response of the price level to an unanticipated disruption in crude-oil supply; cumulative response of the price level to an unanticipated increase in aggregate demand; cumulative response of the price level to an unanticipated increase in oil-specific demand; cumulative response of real GDP to an unanticipated increase in oil-specific demand; cumulative response of real GDP to an unanticipated increase in aggregate demand. Solid lines are based on the point estimates from the model in (4). Shaded areas show one- and two-standard-error bands constructed using a block bootstrap procedure. The estimated structural shocks used to produce the responses are based on data from February 1974 to March 2024.
@@
Turning to the second row of 6, an unanticipated expansion in aggregate demand causes real GDP to increase on impact, although the effect is marginally statistically significant. After 5 quarters, real GDP falls below its initial level. The same unanticipated expansion causes the price level to rise for the first 7 quarters after the shock. The response is statistically significant using the one-standard-error bands. This pattern is consistent with the increase in global aggregate demand causing an initial increase in US growth, but the growth is eventually hampered by the higher price of global commodities, including the price of oil, causing real GDP to fall. Kilian (2009a) provides further discussion.
@@
In summary, we are able to replicate Kilian’s (2009a) substantive results. Patterns uncovered by the historical decomposition and the second-stage impulse response functions hold using data updated through March 2024.
@@
Section 2.4 provides evidence that conclusions about the determinants of the real price of crude oil apply through March 2024; namely, disruptions to oil supply contribute little to fluctuations in price, whereas oil-specific demand shocks are responsible for the ups and downs and aggregate-demand shocks cause “long swings” in price (Kilian, 2009a, 1062). Section 2.5 provides evidence that the US economy responds similarly to changes in the oil market over the later period.
To test this stability proposition further, in this section we use updated data to estimate the impulse response functions. In addition, confidence intervals will be produced using the moving-block-bootstrap method introduced by Brüggemann, Jentsch, and Trenkler (2016). Their recursive-design residual-block-bootstrap procedure is robust to conditional heteroskedasticity. In addition, the method is designed to replicate the asymptotic distribution of both the slope parameters and the variance–covariance matrix. In effect, the wild bootstrap “produces confidence intervals that are too narrow and have coverage accuracy that is too low asymptotically” (Kilian and Lütkepohl, 2017, 346–347). An excellent discussion of the issue is provided by Kilian and Lütkepohl (2017). To implement the moving-block bootstrap procedure, a block length of 36 months is chosen.
Figure 7 shows the responses of production, real economic activity, and the real price of oil to one-standard-deviation structural innovations. When comparing figure 7 to figure 3, there are two differences to keep in mind. First, the structural VAR model is estimated using data through March 2024. Second, the confidence intervals are constructed using the procedure introduced by Brüggemann, Jentsch, and Trenkler (2016). Nevertheless, results are remarkably similar. An unexpected disruption to oil supply causes an immediate decline in oil production and the decline persists. The unexpected fall in oil production initially has no effect on bulk-dry-cargo shipping rates relative to trend, but the measure of real economic activity picks up after 4 months. The rise could reflect higher fuel costs or expected higher fuel costs (relative to the US consumer price index, which is used to deflate the series). For example, the next panel shows that the unexpected disruption to supply causes an even more pronounced rise in the price of oil. The rise in the price of oil is statistically significant at all horizons based on the one-standard-error bands.
@@
Note: Solid lines show the point estimates and the shaded regions show one- and two-standard-error bands. Estimates are based on the structural VAR model in (1), using the updated data (February 1974–March 2024). Confidence intervals were constructed using a residual block bootstrap procedure (Brüggemann, Jentsch, and Trenkler, 2016).
Turning to the second row of figure 7, over the entire sample, oil production seems to respond more acutely to an unanticipated increase in aggregate demand, although a months-long delay remains. The unanticipated rise in aggregate demand causes a large, significant jump in global real economic activity, but the effect is less persistent compared to the effect estimated over only the earlier period. And, compared to only the earlier period, there is a more pronounced rise in the real price of oil in response to an unanticipated increase in aggregate demand. Nevertheless, the patterns are overwhelmingly consistent with Kilian’s (2009a) substantive conclusions.
@@ 4 Application to the clean-energy transition
Much previous interest has centered around the role that oil-supply and -demand shocks have on macroeconomic outcomes like inflation, including responses by the Federal Reserve (see Bernanke, Gertler, and Watson, 1997, for a prominent example). Another compelling reason to study these types of shocks is understanding how local economies depend on oil.
Certain local economies may be especially hard hit as the US transitions away from fossil fuels to sources of clean energy. Previous structural transitions—such as the transition away from coal and the transition to work in the United States being replaced by machines and jobs elsewhere—are causes for concern. In California, policymakers have adopted the goal of carbon neutrality by 2045 or earlier (CARB, 2022). Kern County produces 70 percent of all oil produced in California. As California transitions away from oil, major concerns for Kern include lost jobs and lost tax revenue.
@@
On such a granular level, it is nearly impossible to determine how much of Kern’s economy depends on oil. Yet, one way to understand the linkages between Kern and the global market for oil is by assessing how Kern’s unemployment rate responds to structural innovations identified by the statistical model in (1), using second-stage regressions like those in (4). Specifically, we regress the seasonally adjusted, monthly unemployment rate in Kern on lags of the estimated structural shocks:
$\text{unemployment rate in Kern}_{t}=\kappa+\sum_{i=0}^{12}\psi_{ji}\hat{\varepsilon}_{jt-i},$ (5)
@@
where $\hat{\varepsilon}_{jt}$ refers to the estimated residual in (1) for the $j$th structural shock in month $t$. The impulse response coefficients $\psi_{jh}$ are the response of Kern’s unemployment rate to structural-innovation $j$ at horizon $h$. Inference is based on a block-bootstrap procedure where the block length is chosen to be half a year.
@@
Data on the unemployment rate in Kern County, California, are available from the Local Area Unemployment Statistics program of the Bureau of Labor Statistics starting in January 1990. The data extend through March 2024.
@@
Figure 8 shows the estimated coefficients. The solid, beige lines show the impulse-response coefficients. The shaded areas show the one- and two-standard-error bands.
@@
Figure 8: Responses of the unemployment rate in Kern County, California, to identified structural shocks
@@
Note: Top, response of the unemployment rate to an unanticipated disruption in oil supply; middle, response of the unemployment rate to an unanticipated increase in aggregate demand; bottom, response of the unemployment rate to an unanticipated increase in oil-specific demand. Solid lines show the point estimates from the model in (5). Shaded areas show one- and two-standard-error bands constructed using block bootstrap procedure.
The top panel shows that, upon impact, an unanticipated disruption in oil production has little effect on the unemployment rate in Kern. And there is some evidence that the unemployment rate gradually moves upward, but the evidence is not statistically significant. The response is not obvious, as a disruption to global supply from Kern’s perspective could have represented a windfall increase in the price of oil, which would have benefited the local economy. This does not appear to be the case. Unemployment rises on impact and remains higher a year after the initial disruption, although the effect is marginally statistically significant.
In contrast, an unanticipated one-standard-deviation increase in aggregate demand causes unemployment to decline by over 1 percentage point 4 months after the shock. By 8 months, the unemployment rate falls by 2 percentage points. A year after the shock, the unemployment rate is 3 percentage points lower. The decline is statistically significant at all horizons.
The bottom panel shows that an unanticipated increase in oil-specific demand is also good for the Kern economy. A year after the initial increase the unemployment rate is 2.5 percentage points lower, having fallen throughout the year after the shock. The effect is significant based on the one-standard-error band.
These effects are meaningful. The expanded economic opportunities after an increase in oil-specific demand suggests that Kern may well depend on oil in a meaningful way and the transition away from fossil fuels may be a challenge. Further research on this topic will be aided by having the option to work within the R ecosystem.
@@
Figure 8 makes clear, though, that Kilian’s (2009a) point about identifying the origin of the shock is paramount. Even though the price of oil may be determined independently of economic activity in Kern County, the dynamic effect of a negative supply disruption differs from the dynamic effect of a positive demand innovation—the first causes the unemployment rate to rise, while the second causes the unemployment rate to fall.

### v2

@@ 2.4 Demand and supply shocks in the global oil market affect the US economy in different ways
Figure 4 shows how the level of US real GDP and the level of the CPI respond to the structural shocks. The solid, gray lines show impulse responses estimated using the original data and the replication code. The associated two-standard-error bands are shaded in green. The impulse responses estimated using Kilian’s (2009a) code agree exactly, which produces a striped effect. The dotted, blue lines show the upper and lower ends of the confidence intervals using Kilian’s (2009a) code. These estimates produce slightly narrower confidence intervals, likely due to the fact that the block length used to produce the shaded region is two quarters longer. Nevertheless, the difference is minor. Figure 4 is directly comparable to Kilian’s (2009a) figure 5.
@@
Note: Clockwise from top right, cumulative response of the price level to an unanticipated disruption in crude-oil supply; cumulative response of the price level to an unanticipated increase in aggregate demand; cumulative response of the price level to an unanticipated increase in oil-specific demand; cumulative response of real GDP to an unanticipated increase in oil-specific demand; cumulative response of real GDP to an unanticipated increase in aggregate demand; cumulative response of real GDP to an unanticipated disruption in crude-oil supply. Solid lines are based on the point estimates from the model in (4). Shaded areas show two-standard-error bands constructed using a block bootstrap procedure. Dotted lines show the point estimates and two-standard error bands produced by Kilian’s (2009a) code. All statistics in the figure are based on the original data (Kilian, 2009b).
Figure 5 updates figure 4. To do this, data from February 1974 to January 2025 are first used to estimate the structural VAR model. Then the associated monthly series of structural shocks are averaged by quarter to produce the quarterly averages used in the second-stage regressions in (4).
@@
Note: Clockwise from top right, cumulative response of the price level to an unanticipated disruption in crude-oil supply; cumulative response of the price level to an unanticipated increase in aggregate demand; cumulative response of the price level to an unanticipated increase in oil-specific demand; cumulative response of real GDP to an unanticipated increase in oil-specific demand; cumulative response of real GDP to an unanticipated increase in aggregate demand; cumulative response of real GDP to an unanticipated disruption in crude-oil supply. Solid lines are based on the point estimates from the model in (4). Shaded areas show one- and two-standard-error bands constructed using a block bootstrap procedure. The estimated structural shocks used to produce the responses are based on data from February 1974 to January 2025.
@@
Turning to the second row of figure 5, an unanticipated expansion in aggregate demand causes real GDP to increase on impact, although the effect is marginally statistically significant. After 5 quarters, real GDP falls below its initial level. The same unanticipated expansion causes the price level to rise for the first 7 quarters after the shock. The response is statistically significant using the one-standard-error bands. This pattern is consistent with the increase in global aggregate demand causing an initial increase in US growth, but the growth is eventually hampered by the higher price of global commodities, including the price of oil, causing real GDP to fall. Kilian (2009a) provides further discussion.
@@
In summary, we are able to replicate Kilian’s (2009a) substantive results. Patterns uncovered by the historical decomposition and the second-stage impulse response functions hold using data updated through January 2025.
@@
Section 2.3 provides evidence that conclusions about the determinants of the real price of crude oil apply through January 2025; namely, disruptions to oil supply contribute little to fluctuations in price, whereas oil-specific demand shocks are responsible for the ups and downs and aggregate-demand shocks cause “long swings” in price (Kilian, 2009a, 1062). Section 2.4 provides evidence that the US economy responds similarly to changes in the oil market over the later period.
In this section we use updated data to estimate the impulse response functions. We follow Kilian and Lütkepohl (2017) and construct confidence intervals using the moving-block-bootstrap procedure introduced by Brüggemann, Jentsch, and Trenkler (2016). Kilian and Lütkepohl (2017, pp 415–416) use this procedure to assess Kilian’s (2009a) conclusions. (We are doing nothing new—we are using the same inference procedures, only inference is being carried out over a different sample period.) Brüggemann, Jentsch, and Trenkler’s (2016) recursive-design residual-block-bootstrap procedure is robust to conditional heteroskedasticity. In addition, the procedure is designed to replicate the asymptotic distribution of both the slope parameters and the variance–covariance matrix. In effect, the wild bootstrap “produces confidence intervals that are too narrow and have coverage accuracy that is too low asymptotically” (Kilian and Lütkepohl, 2017, 346–347). An excellent discussion of the issue is provided by Kilian and Lütkepohl (2017). To implement the moving-block bootstrap procedure, a block length of 36 months is chosen.
Figure 6 shows the responses of production, real economic activity, and the real price of oil to one-standard-deviation structural innovations. When comparing figure 6 to figure 2, there are three differences to keep in mind. First, the structural VAR model in (1) is estimated using data retrieved by our code from the sources listed in the text. Second, the data cover the period February 1974 to January 2025. Third, the confidence intervals are constructed using the procedure introduced by Brüggemann, Jentsch, and Trenkler (2016). Nevertheless, results are remarkably similar. An unexpected disruption to oil supply causes an immediate decline in oil production and the decline persists. The unexpected fall in oil production initially has no effect on bulk-dry-cargo shipping rates relative to trend, but this measure of real economic activity picks up after 4 months. The rise could reflect higher fuel costs or expected higher fuel costs (relative to the US consumer price index, which is used to deflate the series). For example, the next panel shows that the unexpected disruption to supply causes an even more pronounced rise in the price of oil. The rise in the price of oil is statistically significant at all horizons based on the one-standard-error bands.
@@
Note: Solid lines show the point estimates and the shaded regions show one- and two-standard-error bands. Estimates are based on the structural VAR model in (1), using the updated data (February 1974–January 2025). Confidence intervals were constructed using a residual block bootstrap procedure (Brüggemann, Jentsch, and Trenkler, 2016).
Turning to the second row of figure 6, over the entire sample, oil production seems to respond more immediately to an unanticipated increase in aggregate demand, although a months-long delay remains. The unanticipated rise in aggregate demand causes a large, significant jump in global real economic activity, but the effect is less persistent compared to the effect estimated over only the earlier period. And, compared to only the earlier period, there is a more pronounced rise in the real price of oil in response to an unanticipated increase in aggregate demand. Nevertheless, the patterns are overwhelmingly consistent with Kilian’s (2009a) substantive conclusions.
@@ 4 Application to the clean-energy transition
Much previous interest has centered around the role that oil-supply and -demand shocks have on macroeconomic outcomes (for example, Kilian, 2008, 2009a, looked at macroeconomic outcomes in response to oil-demand and oil-supply shocks). Another compelling reason to study these types of shocks is understanding how local economies depend on oil.
Certain local economies may be especially hard hit as the US transitions away from fossil fuels to sources of clean energy. Previous structural transitions—such as the transition away from coal and the transition to work in the United States being replaced by machines and jobs elsewhere—are causes for concern. Workers displaced from jobs that are no longer performed are especially hard-hit. And concern extends to the communities in which these workers live (Hanson, 2022). [^4] The forces that restructure the composition of economic activity and the consequences of these forces are not fully understood. Job loss in the coal industry is discussed by Weber (2020). Between 2011 and 2016, the number of people working in mines declined by 45 percent. Each lost coal-mining job is associated with a reduction in county-level income by \$100,000 annually. Retirement of coal-fired power plants in the United States is discussed by Davis, Holladay, and Sims (2022) and the negative effects on workers of shutting down coal-fired power plants are discussed by Hanson (2022); Colmer et al. (2024). Work being automated is discussed by Acemoglu and Autor (2011); Acemoglu and Restrepo (2020, 2022). World trade and China’s participation in world trade is discussed by Autor, Dorn, and Hanson (2021, 2016). The change in composition associated with the hollowing out of middle-pay jobs and when this reallocation takes place over the business cycle is discussed by Autor and Dorn (2013); Foote and Ryan (2014); Howes (2022). Herrendorf, Rogerson, and Valentinyi (2014) provide a framework for structural change within the tradition of economic growth (see also Ngai and Pissarides, 2007). Within this framework, Rogerson (2015) provides a way to think about evaluating environmental regulation.
In California, policymakers have adopted the goal of carbon neutrality by 2045 or earlier (CARB, 2022). Within California, Kern County accounts for 70 percent of all oil produced. As California transitions away from oil, major concerns for Kern include lost jobs and lost tax revenue. Many workers there are primarily engaged in the exploration, development, and production of oil. In addition, the assessed value of oil-field properties fluctuates with factors primarily outside the control of Kern policymakers (Espinoza, 2024). In particular, fluctuations in oil futures prices determine valuations used to assess property taxes. Despite efforts by Kern policymakers to smooth the ups and downs, shocks from the global oil market may propagate to other sectors.
Linkages between labor-market outcomes in oil-producing counties and the oil market have been investigated previously. Michieka and Gearhart (2019, 2021) investigate relationships between nominal regional oil prices and (i) employment and (ii) nominal wages in oil-producing counties. The dynamic-panel models they use are based on the idea that along with the number of time periods getting large, the number of cross-sectional groups is also relatively large. A common “long-run relationship” between regional oil prices and other variables is posited to exist across counties. Importantly, they focused on the causal short- and long-run effects of oil prices (consistent with the econometric specification here that assumes oil prices are exogenous with respect to outcomes in Kern). In contrast, we are interested in the impulse response functions, which trace out the responses of labor-market outcomes in Kern to one-time structural innovations in the global oil market. We are also interested in the labor-market dynamics associated with a disruption to supply and a shift in demand, including whether the shift represents a structural innovation to global real economic activity or precautionary demand. [^5] Michieka and Gearhart (2015) posit a cointegrating relationship between the nominal price of oil and employment in Kern County and estimate a vector error correction model. The rationale for imposing the cointegrating relationship is statistical rather than economic, which is one reason to examine the sensitivity to this modeling assumption. Michieka and Gearhart (2015) do not investigate impulse responses, which is our interest.
@@
It is nearly impossible to definitively state how much of Kern’s economy depends on oil. Yet, one way to understand the linkages between Kern and the global market for oil is by assessing how Kern’s labor market responds to structural innovations identified by the statistical model in (1), using the second-stage methodology proposed by Kilian (2009a). Specifically, we regress the seasonally adjusted labor-market statistics in Kern on lags of the estimated structural shocks:
$k_{t}=\kappa+\sum_{i=0}^{12}\psi_{ji}\hat{\varepsilon}_{jt-i}+e_{jt},$ (5)
@@
where $k_{t}$ is either the unemployment rate or the monthly growth rate of employment, $\hat{\varepsilon}_{jt}$ refers to the estimated residual in (1) for the $j$th structural shock in month $t$, and $e_{jt}$ is a potentially serially correlated error. The coefficient $\psi_{jh}$ is the impulse response at horizon $h$ of series $k$ to structural innovation $j$. Inference is based on a block-bootstrap procedure, where the block length is chosen to be half a year.
@@
Data on the level of employment and the unemployment rate in Kern County, California, are available from the Local Area Unemployment Statistics program of the Bureau of Labor Statistics starting in January 1990. The data extend through January 2025. Both series are retrieved from FRED and seasonally adjusted using the US Census Bureau’s X-13ARIMA-SETS procedure.
@@
Figure 7 shows the estimated coefficients. The solid, beige lines show the impulse-response coefficients. The shaded areas show the one- and two-standard-error bands. Because employment is entered into the statistical model as the change in log employment, we sum the estimated coefficients from (5) to estimate the cumulative response of employment.
@@
Figure 7: Responses of employment and the unemployment rate in Kern County, California, to identified structural shocks
@@
Note: Clockwise from top right, response of the unemployment rate to an unanticipated disruption in oil supply; response of the unemployment rate to an unanticipated increase in aggregate demand; response of the unemployment rate to an unanticipated increase in oil-specific demand; cumulative response of employment to an unanticipated increase in oil-specific demand; cumulative response of employment to an unanticipated increase in aggregate demand; cumulative response of employment to an unanticipated disruption in oil supply. Solid lines are based on the point estimates from the model in (5). Shaded areas show one- and two-standard-error bands constructed using a block bootstrap procedure. The estimated structural shocks used to produce the responses are based on data from February 1974 to January 2025. The regressions that use Kern’s unemployment rate begin in January 1990 and extend through January 2025. The regressions that use the change in employment begin one month later.
The top, left panel shows that, upon impact, an unanticipated disruption in oil production causes employment to rise. This evidence is consistent with the idea that “oil supply contractions in one region tend to trigger production increases elsewhere in the world” (Kilian, 2009a, 1060–1062). The evidence is also consistent with the notion that an increase in the price of oil represents a windfall boost to income from the perspective of some in Kern’s economy. The windfall may propagate throughout the local economy. Employment rises and, turning to the top, right panel, expanded opportunities increase participation in the labor market, as the unemployment rate rises. The effect on the unemployment rate, however, is not statistically significant based on the one- and two-standard-error bands.
In contrast, an unanticipated one-standard-deviation increase in aggregate demand causes the unemployment rate to fall on impact. The effect is statistically significant after 2 months based on the two-standard-error bands. The effect on employment, however, is less pronounced. On impact the effect cannot be distinguished from zero. And while employment expands after two months, the effect is statistically significant based on one-standard-error bands only. These conclusions can be seen in the middle panel of figure 7.
The bottom panel shows that an unanticipated increase in oil-specific demand expands employment opportunities in Kern. On impact, employment rises. And continues to rise until by 7 months the effect is statistically different from zero based on the two-standard-error bands. Employment is higher a year after the initial shock. The right panel shows that the unemployment rate falls on impact. The effect is statistically different from zero after around 9 months. These effects are consistent with the notion that oil production in Kern may respond to oil-specific demand and local revenue from oil production will rise. Linkages between oil extraction and the regional economy propagate to expanded opportunities throughout Kern’s labor market.
The effects suggest that Kern depends on oil in a meaningful way. As the US economy transitions away from dependence on fossil fuels toward green energy, Kern’s economy may be drastically reshaped. What had previously been a major component of regional economic activity may no longer provide opportunities. Policymakers may need to consider expanding options in Kern through, for example, place-based policy. Further research on this topic will be aided by having the option to work within the R ecosystem.
@@
Figure 7 makes clear, though, that Kilian’s (2009a) point about identifying the origin of the shock matters. Even though the price of oil may be determined independently of economic activity in Kern County, the dynamic effect of a negative supply disruption differs from the dynamic effect of a positive demand innovation. A negative supply disruption expands labor-market opportunities more rapidly than the expanded opportunities associated with an oil-specific-demand shock. The effects of a positive demand innovation are visible in the data after 1 year.